OpenStudy (anonymous):

Medal & fan & testimonial for help with this question! Simplify: sqrt75 -4 sqrt8 -3 sqrt32?

OpenStudy (anonymous):

@AaronAndyson @AllanO-the-great @Anaise @Alex_Mattucci @angelnicole @Aileennx3 @bubblysmiles24 @BAdhi @BrandonR @bgrundstrom @babylove2015 @CallMeKiki @Catseyeglint911 @clamin @chainedecho @Directrix

OpenStudy (anonymous):

@zepdrix

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hey, @misty1212 ! Could you help me out?(:

OpenStudy (misty1212):

\[\huge\sqrt{75}-4\sqrt{8}-3\sqrt{32}\]?

OpenStudy (anonymous):

Yes ma'am!

OpenStudy (misty1212):

ok which one you want to do first?

OpenStudy (misty1212):

nvm how about the first one

OpenStudy (anonymous):

Its the whole thing that I have to simplify

OpenStudy (misty1212):

\[\sqrt{75}=\sqrt{25\times 3}\] right?

OpenStudy (anonymous):

Yes

OpenStudy (misty1212):

that means \[\sqrt{75}=\sqrt{25}\times \sqrt{3}=5\times \sqrt{3}\] for the first one

OpenStudy (anonymous):

Ohh!!

OpenStudy (misty1212):

lets move on to the next one

OpenStudy (misty1212):

\[\sqrt{8}=\sqrt{4\times 2}\]yes?

OpenStudy (anonymous):

Yes, so 2 sqrt2?

OpenStudy (misty1212):

bingo

OpenStudy (misty1212):

so far we are at \[5\sqrt{3}-8\sqrt{2}\] the \(-8\) because it started at \(-4\sqrt{8}\)

OpenStudy (anonymous):

Ohh okay

OpenStudy (misty1212):

last one is \(-3\sqrt{32}\) which is the same as \[-3\sqrt{16\times 2}\]

OpenStudy (misty1212):

then since \(\sqrt{16}=4\) you have \[-3\sqrt{16}\sqrt{2}=-3\times 4\times \sqrt{2}=-12\sqrt2\]

OpenStudy (misty1212):

putting it all together (we are still not done) you get \[5\sqrt{3}-8\sqrt2-12\sqrt2\]

OpenStudy (anonymous):

this is called "simplest radical form" btw

OpenStudy (misty1212):

now combine like terms, since the last two both have a \(\sqrt2\) final answer is \[5\sqrt3-20\sqrt2\]

OpenStudy (anonymous):

Okay, so would my final answer be 960 sqrt3

OpenStudy (misty1212):

@satellite73 stop commenting and start helping !

OpenStudy (anonymous):

Whaaat :/

OpenStudy (anonymous):

no that is not the final answer

OpenStudy (anonymous):

\[5\sqrt3-20\sqrt2\] is , as @misty1212 said

OpenStudy (anonymous):

that is because \[-8\sqrt2-12\sqrt2=-20\sqrt2\]

OpenStudy (anonymous):

ohhh we dont multiply.. and why does the 2 stay in the sqrt when there are two of them? Shouldnt it just canel out to be 2?

OpenStudy (anonymous):

the first term has a \(\sqrt3\) in it, you can't combine it with the other two

OpenStudy (anonymous):

good question! because you are adding (subtracting actually) not multiplying

OpenStudy (anonymous):

it is like \[5x-8y-12y=5x-20y\] same idea you aren't multiplying anything, just combining like terms

OpenStudy (anonymous):

Ohhh okay, makes sense! Thanks so much @satellite73 & @misty1212 !

OpenStudy (anonymous):

yw

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]